for all positive integers x, what is the greatest common factor of the 2 numbers 216x and 180x
Answer by greenestamps(13200)



The greatest common factor of 216x and 180x is the greatest common factor of 216 and 180, times the additional common factor x.

216 = 36*6; 180 = 36*5; the GCF of 216 and 180 is 36.

ANSWER: the GCF of 216x and 180x is 36x.